BS03cfg006 = in room 124 and got the BS03+RS04+Splitter01 after the BS+RS+Spltr came out of post-M265 & got pressure washed. Mike and Sean re-terminated the LuBot BS fiber from post-M264 (i.e. 600 um fiber in the black hydraulic tube used during M264), and connected it to port #6 on the BS03+RS04+Splitter01.
RS04cfg003 = in room 124 and got the BS03+RS04+Splitter01 after the BS+RS+Spltr came out of post-M265 & got pressure washed. Mike and Sean re-terminated the LuBot BS fiber from post-M264 (i.e. 600 um fiber in the black hydraulic tube used during M264), and connected it to port #6 on the BS03+RS04+Splitter01.
Day 1 - 8 where collected in Rooom 124. Day 9 data are collected in the tent using the lodestar to make the BS/RS vertical and horizontal
Note: Mike says "I think the BS&RS were left powered on overnight for Art for day6 & 7. The 2x extra fans (blowing across the housings) were added at end-of-day7 after the BS over-heated." This may or may not explain some of the stuff we are seeing in the darks especially for the RS.
Docs which are the reason for this experiment
Andor camera from BS02 - Mike Kehoe fiber bending experiment (PDF)
First-cut at documenting the setup and results for
Mike K.'s "Fiber Optic Stability Experiments: Part 1" (PDF)
Mikes summary document on this bending experiment (PDF)
Problems seen (and solved) and good things:
Log Sheets BS03
Log Sheets RS04 (day01 only, for other says see the BS03 log sheet)
All the raw data files and their images , Table of KEYWORDS variable | |||
Page Number |
Link |
Description |
Date |
1.01 |
DAY01 - first-order check of whether or not the Lu head could be
placed & replaced in front of the sphere repeatably |
10 Jan 2019 |
|
1.02 |
DAY02 - straight FO, 90deg bend near Lu head, straight, 90deg, straight |
11 Jan 2019 |
|
1.03 |
DAY03 - straight FO, 1x FO loop, 2x loops, 1x loop, straight |
12 Jan 2019 |
|
1.04 |
DAY04 - A) compare straight FO vs coiled FO and B) Put a tight-ish loop of FO near the output / spec end |
15 Jan 2019 |
|
1.05 |
DAY05 - A. = straight FO, 90 deg bend @ 24 in dia, straight FO.
B. = straight FO (same straight FO as end of A.), then add
1x 14 in dia loop near spec, then 1x additional 17 in dia loop near Lu head,
then back to only 14 in dia near spec, then back to straight FO. |
16 Jan 2019 |
|
1.06 |
DAY06 - more track 9, via tighter loops @ 11 in outside diameter (noted: the BS+RS were left powered on last night) |
17 Jan 2019 |
|
1.07 |
DAY07 - more track 9, tight-coils:
2x tight loops near the spec & 1 tight 9 deg bend near the Lu head. |
17 Jan 2019 |
|
1.08 |
DAY08 - more track 9, two parts:
A = repeat end-of-day7 with BS & RS;
B = make the FO's tight-90deg-bend at head into a tight-360deg-coil at head,
...then add & remove a 11in dia loop near head. |
22 Jan 2019 |
|
1.09 |
DAY09 - more track 9, vertical/horizontal spec testing, also pretesting images to look at negative data on RS |
6 Feb 2019 |
|
Daily files | |||
2.01 | Daily Track Movement -> BS link & RS link | Track movelength graphs for all days data | 22 Jan 2019 |
2.02 | Repeats and bending -> BS link & RS link | Looking at each days data (repeats, bends and loops) | 22 Jan 2019 |
2.03 | darks-> BS link & RS link | Darks from all experments | 22 Jan 2019 |
2.04 | Low Pixel nets-> BS link & RS link | Mike asked why the RS Low pixel (1-90) data are negative (so I made these graphs) | 22 Jan 2019 |
2.05 | Day 9 trk movment closer look RS link | Closer look at track movement on day 9 | 7 Feb 2019 |
2.06 | Day 9 negative values RS link | Closer look at negative number on day 9 (starting with first data set collected) | 7 Feb 2019 |
DAY 1 EMAIL from Mike |
Hi Steph! How are you doing?
We have some more BS+RS data for thee! I think Mark brought this up during the last telecon (?) - the idea of checking the effects of bending a fiber optic using the Refresh 600 um fiber + collett fiber termination + collimators. Last week I cleared out space in room 124 and got the BS03+RS04+Splitter01 in there. I.E. after the BS+RS+Spltr came out of post-M265 & got pressure washed. This week Sean & I re-terminated the LuBot BS fiber from post-M264 (i.e. 600 um fiber in the black hydraulic tube used during M264), and connected it to port #6 on the BS03+RS04+Splitter01. We also dis/re-connected the FO at the LuBot head to verify its' veracity. Today we took a first set of data. The light source was the NIST OL455-18U, with photodiode monitor Amp output logged. The scan data were all sets of 2dk/4lt/2dk, at 5 sec exp time, at -60 degC TEC. Data were collected via asg_interactive software, to s* & d*.fits files - i.e. sig & drk scans using the "external" BS & RS shutters. Tonight was a first-order check of whether or not the Lu head could be placed & replaced in front of the sphere repeatably. Ex. take 1 scan set, then take 4x more scan sets to look at stability, then remove the Lu head from the mount @ sphere & replace, and take 4x more scan sets to look at stability, then repeat the head remove/replacement 2x more times. I called this experiment HI-2019-01. Tonight was day01. I am VPN-ing data now to your NAS at Mike/HI-2019-01/. There are BS03 & RS04 .fits files + 2x putty logs + 2x log sheets, and setup photos. There is 1x OL455 PD monitor log in the BS3cal/ directory. If these data make sense to you, then the hope/plan is to try tomorrow to put some bends into & remove from the FO & move the sphere & re-scan. Mark's intent is to see if a couple of tight FO coils near the BS/RS output-end of the FO will remove sensitivity to FO bending between the tight coils and the input collector - like Carol's off-the-shelf instrument. Make sense? The VPN kine from home tonight seems to be slow-butt-steady: copying from hard disc to/from USB memory stick goes at 8 or 10 M bytes/sec, copying PC hard disc to NAS via VPN goes at ~190 K bytes/sec. Aloha, MF |
DAY 2 EMAIL from Mike |
HI StePh, Day02 data are slowly heading to the NAS. Tonights' data were: straight FO, 90deg bend near Lu head, straight, 90deg, straight. I used only 1 log sheet for both the BS & RS (and had to rename the day01 log sheets, sorry). There are new day02 photos + 2x putty logs + 1x OL455 PD log file. I'm very curious if the top-hat track-shape from day01 straight FO becomes saddle-shaped with the day02 90deg FO bend. And I'm maybe more curious if the signal repeats when the fiber is un-bent or re-bent! At Mark's suggestion we added a T-Slot rail to help keep the straight portion of the FO straight. And we added a pair of fans blowing on the BS & RS radiator fins. The plan for day03 is to put a loop or two in the FO near the output end and see if the coils reduce sensitivity to FO bending between the coils and the input end @ collector head. The diameter of the coils may be impotent / need reducing... MF |
DAY 3 EMAIL from Mike |
Hi Steph, I am VPNing day 3 data to your NAS: 1x log sheet, 1x OL455 log, 2x putty logs, 2x 163x .fits files, 11x day3 photos. The measurements sequence was: straight FO, 1x FO loop, 2x loops, 1x loop, straight. I tried to get the repeats of straight & 1x loop in aboot the same physical place. The FO loops were not tight @ ~18 in diameter. I'm guessing the next step might be tighter loops... Should be interesting what you see! MF |
DAY 4 EMAIL from Mike |
Hi Steph, Day4 data are heading to the NAS. There were 2 sub-experiments today: A. = compare straight FO vs coiled FO, similar to how we had originally thought calibrations could be done (coiled) before shipping out to be installed in a buoy (uncoiled). For this bottom-length FO there were ~4x loops on a FO spool of ~14 inch diameter. B. = Put a tight-ish loop of FO near the output / spec end, ~14 in dia, and see if that decreases the sensitivity of the remaining FO to having another loop closer to the input / head end. There are 1x log sheet + photos + OL455 & putty logs + .fits files at the kine. The last scan set for the BS had some zero-byte files because the BS disc filled up. I reckon some of this should be repeated on another track tomorrow. I know Mark would like to see TIGHT loops of FO, but I'm afraid of breaking / damaging said FO, so I'd like to leave that until the end. Luck, MF |
Email about addition ideas for tests |
Hi Steph, Mark came in today and re-stated his theory that we need TIGHT loops to de-sensitize the FO to bending, and by TIGHT he thinks 2-3 in dia loops. This will ruin the fiber to further use. M also thinks the TIGHT loops need to be combined with a TIGHT right-angle bend at the input-end of the FO. I.E. like Carol's spec+fiber+head setup. I'm thinking that today = day5 will be trying another track - e.x. T13, with some repeat of the day2-3 scheme. I'm encouraged by the day4 part B data. It looks like a single loop near the spec, may have led to little/no change via adding a second loop near the head... MF |
DAY 5 EMAIL from Mike (day 5 not day 6) |
Hi Steph,
Day6 data are on da NAS. Day6 data were via Trk#9, and they had two parts: A. = straight FO, 90 deg bend @ 24 in dia, straight FO. B. = straight FO (same straight FO as end of A.), then add 1x 14 in dia loop near spec, then 1x additional 17 in dia loop near Lu head, then back to only 14 in dia near spec, then back to straight FO. There are 1x log sheet + 1x OL455 log + 2x putty logs + photos & .fits files. Here's one interesting thing to add to your notes for M265: the collimator in the FO block for Trk#9 had loose lenses. This collimator is at the output end of the Fiber Optic. There are two lenses involved, held in place inside a metal tube via a locking ring. Then this locking ring is locked in place by a second ring. The second ring was tight, but the ring against the lens was not tight, so the lenses were loose. Not terribly loose, but loose enough that I felt them jiggle when I handled the collimator tube. I had to take off said tube from port #9 to install the FO @ port #9, which is when I felt the jiggling lenses. Cheers, MF |
DAY 6 EMAIL from Mike |
Hi Steph, The real day6 data are on the NAS now. Same track#9 input, via tighter loops @ 11 in outside diameter. 1x log sheet, 1x OL455 log, 2x putty logs, 4x photos, whole bunch of .fits files. May it be noted: the BS+RS were left powered on last night (Tuesday HST, after day5 data collection) so Art could work with them this morning (Wednesday HST), then Sean & I ran the BS+RS for day6 data - so, the internal guts of the BS&RS might have been (or might not have been) warmer today than on previous days. Also, The BS+RS were left powered on again tonight (after day6 data collection) so Art can work with them tomorrow morning (Thursday HST). Thanks for keeping up with all of this, MF |
DAY 7 EMAIL from Mike |
Hi Steph, Day7 data are heading toward the NAS. These were tight-coil data. I.E. 2x tight loops near the spec & 1 tight 9 deg bend near the Lu head. With 1x log sheet & OL455 log + photos & blah blah blah. One problem: near the end of the run the BS cooler switched off, presumably because the camera was too hot - a too-hot camera sounds a continuous beep, which could be faintly heard through the plastic housing. So, we had to turn off the BS and finish with only the RS. Which means there are missing files from the BS. And, the last N .fits files from the BS were not at -60 degC - from watching max ADU printout on the screen I think it was only the last scan set from the BS that was bogus, butt, it will be interesting to see from your data listing exactly when the BS CCD temp went above -60 degC. I suppose we'll want to repeat the end of day7 tomorrow, day8 with both the BS & RS... MF |
DAY 8 EMAIL from Mike |
Hi Steph, Day8 data are now on the NAS. Day8 had two parts: A = repeat end-of-day7 with BS & RS; B = make the FO's tight-90deg-bend at head into a tight-360deg-coil at head, ...then add & remove a 11in dia loop near head. I'm feeling like it is time to get back to MOBY267 pre-cals in the tent... Two "issues" there were with data on the NAS: 1. I messed up the file name index at the start of day8 data acquisition. The first set of scan-sets were sequence 001 to 004 - not 001 to 005 as were day1 : day7. Sorry. 2. I messed up the file naming for photos from aboot day5 to day6. The filenames on the NAS are correct now (I hope). I say "aboot" because I had to re- re-name a bunch of photos on the NAS and I am no longer certain which ones really got re-named. Sorry. Luck, MF |
DAY 9 EMAIL from Mike |
HI Stephanie, Do you have time to look at some more BSRS data today? If so, then there are day9 data on the NAS = 1 log sheet + 1 OL455 log + 2x putty logs + 4 photos + 240*2*2 .fits files for the tipping of the specs on-end four times. Let me know if I forgot to copy some data! Thank you, MF |